# Learning thermodynamically constrained equations of state with uncertainty

### v1

@@ IV.3 Unified GP EOS for Diamond Trained from Simulations and Experiments
In this section, we present a unified physics-informed GP EOS for the diamond phase of carbon that is trained from multiple data sources and provides accurate EOS predictions with uncertainty – as described in Section III.3. The given DFT-MD simulation data (same 20 data points used above) is augmented by Hugoniot experimental data from dynamic shock wave experiments that relate pressure and volumeMcWilliams et al. (2010) (3 points). These data are from a pressure regime where the diamond has significant strength, meaning that we expect to see a discrepancy between simulation and experiment which provides an interesting test for our unified modeling approach. Additionally, the temperature is not usually measured in these high-pressure experiments (except in those with static compression), which provides another difficulty to overcome with our approach. The GP EOS model trained in the previous sections can be used to estimate temperature for a given volume using the mapping ($V_{H}\rightarrow T_{H}$) shown in Figure 4(b). We apply a deterministic mapping here, but theoretically, we could also construct a GP relating temperature and volume, which would formally provide a probability distribution for temperature. However, such an approach requires a more challenging deep GP framework (since the input state $T$ is a GP), which is beyond the scope of the present work. With these temperature estimates, we now have a training set $\{\mathbf{X}_{P},P$} (23 data points) and $\{\mathbf{X}_{E},E\}$ (20 data points). We use this training data to train the unified GP. The results are shown in Figure 6. As expected, the uncertainty reduces significantly near the experimental observations in Figure 6(a). Comparing Figure 6(a) with Figure 2(a), we can observe an increase in the uncertainty, which captures the inconsistencies between the simulation and experimental data. Since the P-GP and E-GP are correlated, a slight increase in uncertainty is also observed in Figure 6(b) compared to 2(b). Finally, as shown in Figure 7, the unified GP EOS still satisfies the thermodynamic stability constraints as necessary. We have therefore developed a comprehensive unified framework that has been successfully trained using both first-principles simulation data and experimental shock compression experiments.
@@
Figure 7: Plots of the thermodynamic stability conditions at each point on a $50\times 50$ grid of $(V,T)$ for the unified constrained GP EOS trained on simulation and experimental data. Regions of constraint violation are shaded. (a) Probabilistic stability constraint for unified constrained Pressure GP EOS model showing that the constraints are satisfied. (b) Probabilistic stability constraint for unified constrained Energy GP EOS model showing that the constraints are satisfied.
@@ V Conclusion
The proposed framework can be similarly applied to different material phases. However, an extension of the proposed framework to a more generalized multiphase EOS model that captures phase transitions is the subject of further work. Finally, we anticipate that the proposed framework opens the door to a wide range of improvements in EOS modeling and the associated downstream applications. For example, in future studies the prediction uncertainty can be used to inform the choice of points for new simulations/experiments through e.g. Bayesian optimization resulting in smaller data sets requirements and thus accelerating the development of new EOSs. Moreover, the proposed framework can be integrated with physics-based parametric models such that the physically constrained GP serves as a correction to the parametric model and potentially facilitates multi-fidelity modeling. Finally, these physics-informed GP EOS models will be integrated into hydrocode simulations of shock experiments to enable uncertainty quantification studies.
@@
Code developed in this work and presented results will be made publicly available via GitHub upon publication.

### v2

@@ IV.3 Unified GP EOS for Diamond Trained from Simulations and Experiments
In this section, we present a unified physics-informed GP EOS for the diamond phase of carbon that is trained from multiple data sources and provides accurate EOS predictions with uncertainty – as described in Section III.3. The given DFT-MD simulation data (same 20 data points used above) is augmented by Hugoniot experimental data from dynamic shock wave experiments that relate pressure and volumeMcWilliams et al. (2010) (3 points). These data are from a pressure regime where the diamond has significant strength, meaning that we expect to see a discrepancy between simulation and experiment which provides an interesting test for our unified modeling approach. Additionally, the temperature is not usually measured in these high-pressure experiments (except in those with static compression), which provides another difficulty to overcome with our approach. The GP EOS model trained in the previous sections can be used to estimate temperature for a given volume using the mapping ($V_{H}\rightarrow T_{H}$) shown in Figure 5(b). We apply a deterministic mapping here, but theoretically, we could also construct a GP relating temperature and volume, which would formally provide a probability distribution for temperature. However, such an approach requires a more challenging deep GP framework (since the input state $T$ is a GP), which is beyond the scope of the present work. With these temperature estimates, we now have a training set $\{\mathbf{X}_{P},P$} (23 data points) and $\{\mathbf{X}_{E},E\}$ (20 data points). We use this training data to train the unified GP. The results are shown in Figure 7. As expected, the uncertainty reduces significantly near the experimental observations in Figure 7(a). Comparing Figure 7(a) with Figure 2(a), we can observe an increase in the uncertainty, which captures the inconsistencies between the simulation and experimental data. Since the P-GP and E-GP are correlated, a slight increase in uncertainty is also observed in Figure 7(b) compared to 2(b). Nonetheless, the overall uncertainty is still relatively low in terms of COV. However, if the data is highly inconsistent or a different interpretation of the difference between model and experiment is desired, various approaches from the literature can be considered to address the inconsistency. For example specific analysis may choose a suitable kernel (e.g., Matérn kernel with higher smoothness parameters), more robust likelihood functions (e.g., Huber likelihood), regularizations, and cross-validations among othersWilliams and Rasmussen (2006), while still ensuring physical consistency in the manner we have described. Finally, as shown in Figure 8, the unified GP EOS still satisfies the thermodynamic stability constraints as necessary. We have therefore developed a comprehensive unified framework that has been successfully trained using both first-principles simulation data and experimental shock compression experiments.
@@
Figure 8: Plots of the thermodynamic stability conditions in the $(V,T)$ space for the unified constrained GP EOS trained on simulation and experimental data. Regions of constraint violation are represented by red. (a) Probabilistic stability constraint for unified constrained Pressure GP EOS model showing that the constraints are satisfied. (b) Probabilistic stability constraint for unified constrained Energy GP EOS model showing that the constraints are satisfied.
@@ V Conclusion
The proposed framework can be similarly applied to different material phases. However, an extension of the proposed framework to a more generalized multiphase EOS model that captures phase transitions is the subject of further work. Finally, we anticipate that the proposed framework opens the door to a wide range of improvements in EOS modeling and the associated downstream applications. For example, in future studies the prediction uncertainty can be used to inform the choice of points for new simulations/experiments through e.g. Bayesian optimization resulting in smaller data sets requirements and thus accelerating the development of new EOSs. Moreover, the proposed framework can be integrated with physics-based parametric models similar to the KOH approachKennedy and O’Hagan (2001), such that the physically constrained GP serves as a correction to the parametric model and potentially facilitates multi-fidelity modeling. Finally, these physics-informed GP EOS models will be integrated into hydrocode simulations of shock experiments to enable uncertainty quantification studies.
@@
Code developed in this work, training data, and presented results are available through the following Github repository: https://github.com/himanshu258/constrained_GP_EOS Sharma (2023).
@@ Acknowledgements.
APPENDIX
The Hugoniot equation can be expressed as,
$H(V,T)=E(V,T)-E_{0}+\frac{1}{2}\left(V-V_{0}\right)\left(P(V,T)+P_{0}\right)$ (A.1)
$H=f\left(P,E\right)=f\left(\mu_{P},\mu_{E}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial P}\right|_{\mu_{P},\mu_{E}}\left(P-\mu_{P}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial E}\right|_{\mu_{P},\mu_{E}}\left(E-\mu_{E}\right)$ (A.2)
Since both $P$ and $E$ are GPs, $H$ (a linear function of $P$ and $E$) is also a GP. Taking the expectation of Eq. (A.2) yields:
$\mathop{{}\mathbb{E}}\left[H\right]=\mathop{{}\mathbb{E}}\left[f\left(P,E\right)\right]=f\left(\mu_{P},\mu_{E}\right)=\mu_{H}$ (A.3)
The covariance of the $H$ GP is obtained by:
$\mathop{{}\mathbb{E}}\Big{[}(H-\mathop{{}\mathbb{E}}[H])(H-\mathop{{}\mathbb{E}}[H])\Big{]}$ $=\mathop{{}\mathbb{E}}\bigg{[}\bigg{(}\left.\frac{\partial f\left(P,E\right)}{\partial P}\right|_{\mu_{P},\mu_{E}}\left(P-\mu_{P}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial E}\right|_{\mu_{P},\mu_{E}}\left(E-\mu_{E}\right)\bigg{)}$ (A.4)
The cross-covariance of the $H$ and $P$ GPs is given by
$\mathop{{}\mathbb{E}}\Big{[}(H-\mathop{{}\mathbb{E}}[H])(P-\mathop{{}\mathbb{E}}[P])\Big{]}$ $=\mathop{{}\mathbb{E}}\bigg{[}\bigg{(}\left.\frac{\partial f\left(P,E\right)}{\partial P}\right|_{\mu_{P},\mu_{E}}\left(P-\mu_{P}\right)+\left.\frac{\partial f\left(P,E\right)}{\partial E}\right|_{\mu_{P},\mu_{E}}\left(E-\mu_{E}\right)\bigg{)}\bigg{(}\left(P-\mu_{P}\right)\bigg{)}\bigg{]}$ (A.5)
$K_{HE}\left(\mathbf{X},\mathbf{X}^{\prime}\right)=K_{EE}\left(\mathbf{X},\mathbf{X}^{\prime}\right)+K_{EP}\left(\mathbf{X},\mathbf{X}^{\prime}\right)$ (A.6)
For cases where the GP EOS is intended for applications where a linear relationship may not exist, we can still derive the mean and covariance model in the same manner by considering higher-order closure terms in the Taylor series approximation. Constantinescu and Anitescu (2013) demonstrated how to address non-linear dependencies between GP outputs for physics-based systems to obtain a valid covariances model. There are also recent works in the multioutput GP literature that deals with obtaining covariance models for systems that are non-linear while ensuring positive definitenessHu and Wang (2021); Álvarez, Ward, and Guarnizo (2019).